OpenStudy (anonymous):

Fairview High School has a total of 92 boys and girls who play sports. If the number of boys is 17 more than twice the number of girls, how many boys play sports at this high school?

terenzreignz (terenzreignz):

First step to solving word problems like these is assigning values to stuff.. or what we call representing. If we let x be the number of boys... how would you relate that to the number of girls?

OpenStudy (anonymous):

y equals the number of girls?

terenzreignz (terenzreignz):

Ok... you can do that :) Have you done systems of linear equations already?

OpenStudy (anonymous):

some but im having trouble

terenzreignz (terenzreignz):

Okay... more on that later :) So, x boys, y girls... catch me so far?

OpenStudy (anonymous):

yess

terenzreignz (terenzreignz):

First... what's x + y?

OpenStudy (anonymous):

x+y=92

terenzreignz (terenzreignz):

Very good :) We have two variables, so we need two equations to be able to solve it.. Any idea where we can get another equation?

OpenStudy (anonymous):

17+2y=x?

terenzreignz (terenzreignz):

...Yeah actually... are you sure you needed help with this question? ;)

OpenStudy (anonymous):

lol i doubt myself just making sure

terenzreignz (terenzreignz):

Okay... so... we have x + y = 92 and 17 + 2y = x We want the number of boys, right? So we're looking for x.

OpenStudy (anonymous):

im good now as long as i know that those are the two equations thank you :)

terenzreignz (terenzreignz):

Well, I'll be :) You pretty much solved this by yourself, so... take this medal, and may you have more to come :D By the way, you're new to Openstudy. I may not be an Ambassador, but welcome to Openstudy... and enjoy! :)